OpenStudy (anonymous):

find the slope of the line that passes through (3,-4) and (-1,2)

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

m = (y2 -y1)/(x2-x1) , -4-2 / 3-(-1) = -3/2